Suppose that you buy a share of stock for 28 dollars, you sell one call for 2 dollars, and you buy three puts for 7 dollars each. Assume that all the options have a strike price of 46 dollars, and the same expiration date. If the stock price on the expiration date is 41 dollars, what is your total profit?
Answer #1

Stock price on expiration is 41


So, profit on share = 41 – 28 = 13

Profit on call = 2 - max (41 – 46, 0) = 2

Profit on one put = max (46-41, 0) – 7 = -2

Profit on 3 puts=3*(-2) = -6

So, total profit= 13 + 2 - 6 = 9